Find the slope-intercept form of the line that contains the following points: (–6, –6) and (–3, 1).
Answer by checkley71(8403)

(1+6)/(-3+6)
7/3 is the slope
now substitute one of the point x & y values & solve for the y intercept (b)
1=7/3*-3+b
1=-7+b
b=1+7
b=8 thus the equation of this line is
y=7/3x=8